# On the feasibility of performing quantum chemistry calculations on quantum computers

### v3

@@ I.2.3 Example of the depolarizing noise
To understand better the behaviour of $E_{\infty}$, we perform another study: a linear chain of $N$ hydrogen atoms (a popular benchmark). The left panel of Figure 3 shows the energy $E_{\infty}/N$ versus $N$ in different basis sets, calculated with the PySCF package [32]. The figure confirms the above points (i) and (ii). Also shown is the Hartree-Fock energy (empty symbols, almost basis set independent at this scale) and the maximum energy of the Hamiltonian ($E_{2^{n}}$, star, only for STO-3G) which is, as expected, of the same order of magnitude. However in this figure, point (iii) is not visible: after an initial rise, $E_{\infty}/N$ actually saturates, there is no building up of a macroscopic dipole. This is due to the high symmetry of the problem: except on the two ends, all the atoms are essentially equivalent, hence the electrons are distributed evenly by this noise model. In real applications, however, this is seldom the case as molecules are made of different atoms with different environments. To showcase a more generic case, we have modified the basis set of the hydrogen atom chain in the following way: the first $N/2$ atoms are treated with a small basis set (STO-3G) while the last $N/2$ atoms are treated with a larger basis set. This, arguably artificial, setting should be understood as a pedagogical extreme limit. The results are shown in the right panel of Figure 3. We observe a strong $N$ dependence (iii) with $E_{\infty}$ at least ten times larger than in the symmetric setting. In the simple error model considered here, ($\rho_{\rm noise}\propto I_{\rm d}$), the noise essentially populates the available orbitals in a uniform way. The point we are trying to make is that such a population is likely not to properly screen the nuclei charge, therefore have a dire cost in energy. This is shown in a particularly striking way on the right panel, where the asymmetry in the basis set leads to the advent of a macroscopic dipole (and the energy increases quadratically with $N$: point (iii)). Even for the case on the left, which is highly symmetric (all atoms play an identical role except for the two on the edges of the chain), the different orbitals are not identical, so that increasing the basis set is very costly in energy (point (ii)). In anycase, the most important information from these curves is the scale of the noise energy: of the order of $10$ Ha, i.e. four orders of magnitude larger than the targeted accuracy (point (i)).
@@
Figure 3: Different energy scales of a chain of $N$ hydrogen atoms. (Left) symmetric setting: all atoms are treated with the same basis set. Shown are the Hartree-Fock energy $E_{HF}$ and infinite temperature energy $E_{\infty}$ per atom, for the ”STO-3G” ($2$ qubits per atom), ”6-31G” ($4$ qubits per atom)  [33], ”cc-pVDZ” ($10$ qubits per atom) and ”cc-pVTZ” basis ($28$ qubits per atom)  [34]. For the ”STO-3G” basis we have also plotted the highest excited state energy $E_{max}$ up to $N=10$ (regime where exact diagonalization is possible). In all curves, the energy per atom is offset by its Hartree-Fock value for $N=2$ in the ”STO-3G” basis. (Right) assymetric setting: for the rightmost-half of the chain we use the ’sto-3g’ basis while for the leftmost-half we use more precise basis sets: ”cc-pvdz”, ”dzp”, ”cc-pvtz”, ”tzp”.
I.2.4 Application: criterion estimation for Benzene
@@
We are now in position to answer the question: what does the criterion Eq. (8) imply in practice? It is generally accepted that a quantum chemistry calculation must reach ”chemical accuracy”, ${\eta_{\rm chem}=1{\rm kcal/mol}\ \approx 1.6{\rm mHa}\approx 500K}$ (for energy differences). As we have seen, a very conservative estimate for the noise energy is $E_{\rm noise}\geq 1{\rm Ha}$, likely much more. For the $H_{2}$ molecule in a minimum basis set of just two orbitals per atom (STO-3G), one gets ${E_{\infty}-E_{0}=1.02{\rm Ha}}$ for the depolarizing noise, so that we estimate ${E_{noise}-E_{V}}$ to be of the order of $1$Ha in this simple case. On the other hand, one needs a variational ansatz expressive enough to reach $\eta_{\rm chem}$, which provides a constraint $N_{g}\geq N_{g}(\eta_{\rm chem},N)$.
@@
To be concrete, we consider a recent blind test benchmark on the benzene molecule [35]. Benzene is a nontrivial calculation for classical approaches. Yet [35] showcased that a variety of classical techniques arrived at chemical precision using $30$ electrons distributed on $108$ orbitals. It can be considered as a minimum target for VQE: anything smaller is likely to be addressable by classical means. Using the UCC ansatz, inspired by the successful coupled cluster approach used in quantum chemistry, would require to include at least single, double and triple excitations (actually quadruple would probably be needed too, again we are being optimistic), which translates into $N_{g}>N^{6}$ gates. One arrives at a noise level $\epsilon\leq 10^{-3}/(30)^{6}$, which translates into $\epsilon\leq 10^{-12}$, that is many orders of magnitude below the best existing quantum hardware.
@@
Such a low level of noise is difficult to imagine, even with error mitigation (which we briefly discuss below), without a full-fledged implementation of quantum error correction. Since quantum error correction itself comes with an important overhead in term of computational time, one wants to avoid the accumulation of statistics required in VQE. We will thus turn to a different algorithm, QPE, in section II.
I.2.5 Prospects of error mitigation
As we have observed in the previous section, decoherence is harmful because it introduces a large bias to the sought-after energy. This reason for this bias being large is twofold: noise generically produces high-energy states, and does so exponentially quickly.
@@ I.2.5 Prospects of error mitigation
To fight this deleterious influence, error mitigation techniques have been developed over the last few years [36, 37, 38, 39, 40, 41]. These techniques, including probabilistic error cancellation, zero-noise extrapolation, assignment error mitigation, etc., essentially consist in reducing the bias on the energy estimator using more or less precise information on the noise processes at play. This comes at a steep price: the variance of the energy estimator increases exponentially with the number of gates of the circuit, and even, in some cases, the number of qubits [42, 43, 44]. This in turns leads to a number of required shots, and thus a run time, that increase exponentially. We therefore think it is unlikely that error mitigation will solve the noise problem discussed above: in practice, the number of shots cannot be scaled exponentially with system size, leading to a residual bias. In any case, the above criterion still holds in the presence of practical error mitigation: One just needs to replace $\epsilon$ by $\epsilon^{\mathrm{\rm mitigated}}$ in Eq. (8). For small molecules with less than $15$ orbitals, it has been estimated [45] that error mitigation might improve by one or two orders of magnitude the maximum allowed error per gate for VQE. For larger molecules, because of the scaling problems mentioned above, the gain one can expect from mitigation will be even lower, leaving the threshold of $10^{-12}$ for Benzene out of reach.
I.3 VQE versus its classical competitor, the variational Monte-Carlo (VMC) algorithm
Comparison between quantum and classical approaches
@@ I.3 VQE versus its classical competitor, the variational Monte-Carlo (VMC) algorithm
(III) Imprecisions & decoherence Need ”fault-tolerant quantum computer” level of precision (this work) Works on existing CPU, GPU and TPU. Trivial parallelism.
@@
Table 1: VQE versus VMC for the three main difficulties faced by these algorithms.
In the previous section, we discussed a necessary condition for VQE to be able to reach chemical accuracy. Here, we stress the fact that a condition for quantum advantage is more stringent: VQE must be faster than competing purely classical algorithms. A natural comparison point for VQE is the classical algorithm of variational Monte-Carlo (VMC) [46]—an old technique that has been experiencing an important revival in the last few years thanks to the success of deep neural networks based ansatz. Indeed, the introduction of neural network variational wave functions with a large number of parameters $N_{\theta}$ has led to very encouraging results [47, 48]. In VMC, one also considers a variational wavefunction $|\Psi_{V}(\vec{\theta})\rangle$. This function can be anything as long as one is able to compute its value $\langle\vec{s}|\Psi_{V}\rangle$ for a given configuration $|\vec{s}\rangle$. This includes wavefunctions made of neural networks as well as physically-motivated wavefunctions, or combinations thereof. In VMC, the variational energy is expressed as
@@
The sum over configurations in (14) is sampled by a Monte-Carlo algorithm, namely one draws $N_{s}$ samples from the distribution $|\langle\vec{s}|\Psi_{V}\rangle|^{2}$ and estimates the energy $E_{V}$ as
@@
Here, even if $|\Psi_{V}\rangle$ happens to be the exact ground state, it is not an eigenstate of the individual $P_{\alpha}$ operators and the variance thus does not vanish. In fact, since those terms originally stem from the kinetic and Coulomb repulsion matrix elements, they have a quite large magnitude (typically in the $1$ Ha range). This absence of variance reduction in VQE reflects the fact that it is much more advantageous to measure the full Hamiltonian at once than its different components separately. In practice, as the optimization proceeds in VMC, one wants to measure $E_{V}$ with more and more digits of precision. Since $|\Psi_{V}\rangle$ gets closer to $|\Psi_{0}\rangle$, in VMC the energy variance decreases and hence the precision of the calculation also becomes better for the same experimental time (same number of samples). In contrast, in VQE the variance of the different energies $E_{\alpha}$ does not decrease, hence one must increase the number of samples by a factor 100 for each new digit in precision on the measure of $E_{V}$. This might constitute another important advantage in favor of VMC.
@@
(5) The two techniques have access to different sorts of ansätze. VMC is more versatile in that respect and can simulate VQE ansätze (as long as the quantum circuit depth is not too deep) while the reverse is not true. For instance, VMC can use Tensor network states that give an efficient representation of weakly entangled states. But VMC can also use neural network states which can sustain highly entangled states including volume law entanglement. In VQE, the quantum nature of the hardware allows in principle to build arbitrarily entangled states, but in practice the depth might be excessive. This is particularly true for the existing (NISQ) hardware which is limited to shallow circuits, hence limited entanglement. At the moment it is unclear if one technique has an advantage in that respect. One should keep in mind that the expressivity of the ansatz is strongly linked to the ease/difficulty to perform the optimization. An ansatz which is too expressive (for instance the so-called ”hardware efficient ansatz” in VQE) leads to barren plateaus as most of the parameter space corresponds to almost random (chaotic) states (see the above discussion on barren plateaus).
The conclusions of this comparison between VQE and VMC are summarized in Table 1. What one should take away from this exercise is that, in contrast to e.g. Shor’s algorithm, which is expected to speed up the factorization of large numbers exponentially, there are no similar indications for VQE, quite the opposite actually. Even if we had perfect quantum computers, VMC might still remain superior to VQE. (See also [50] for a discussion including other classical Monte-Carlo methods).

### v4

@@ I.2.3 Example of the depolarizing noise
To understand better the behavior of $E_{\infty}$, we perform another study: a linear chain of $N$ hydrogen atoms (a popular benchmark). The left panel of Figure 3 shows the energy $E_{\infty}/N$ versus $N$ in different basis sets, calculated with the PySCF package [35]. The figure confirms the above points (i) and (ii). Also shown is the Hartree-Fock energy (empty symbols, almost basis set independent at this scale) and the maximum energy of the Hamiltonian ($E_{2^{n}}$, star, only for STO-3G) which is, as expected, of the same order of magnitude. However in this figure, point (iii) is not visible: after an initial rise, $E_{\infty}/N$ actually saturates, there is no building up of a macroscopic dipole. This is due to the high symmetry of the problem: except on the two ends, all the atoms are essentially equivalent, hence the electrons are distributed evenly by this noise model. In real applications, however, this is seldom the case as molecules are made of different atoms with different environments. To showcase a more generic case, we have modified the basis set of the hydrogen atom chain in the following way: the first $N/2$ atoms are treated with a small basis set (STO-3G) while the last $N/2$ atoms are treated with a larger basis set. This, arguably artificial, setting should be understood as a pedagogical extreme limit. The results are shown in the right panel of Figure 3. We observe a strong $N$ dependence (iii) with $E_{\infty}$ at least ten times larger than in the symmetric setting. In the simple error model considered here, ($\rho_{\rm noise}\propto I_{\rm d}$), the noise essentially populates the available orbitals in a uniform way. The point we are trying to make is that such a population is likely not to properly screen the nuclei charge, and therefore to have a dire cost in energy. This is shown in a particularly striking way on the right panel, where the asymmetry in the basis set leads to the advent of a macroscopic dipole [and the energy increases quadratically with $N$: point (iii)]. Even for the case on the left, which is highly symmetric (all atoms play an identical role except for the two on the edges of the chain), the different orbitals are not identical, so that increasing the basis set is very costly in energy [point (ii)]. In anycase, the most important information from these curves is the scale of the noise energy: of the order of $10$ Ha, i.e. four orders of magnitude larger than the targeted accuracy [point (i)].
@@
Figure 3: Different energy scales of a chain of $N$ hydrogen atoms. Left panel: symmetric setting: all atoms are treated with the same basis set. Shown are the Hartree-Fock energy $E_{HF}$ and infinite temperature energy $E_{\infty}$ per atom, for the ”STO-3G” (22 qubits per atom), ”6-31G” (44 qubits per atom)  [36], ”$c$-pVDZ” ($10$ qubits per atom) and ”$c$-pVTZ” basis ($28$ qubits per atom)  [37]. For the ”STO-3G” basis we have also plotted the highest excited state energy $E_{max}$ up to $N=10$ (regime where exact diagonalization is possible). In all curves, the energy per atom is offset by its Hartree-Fock value for $N=2$ in the ”STO-3G” basis. (Right) Effect of an assymetry on the large energy spectrum. For the rightmost-half of the chain we use the ’sto-3g’ basis while for the leftmost-half we use more precise basis sets: ”$c$-pvdz”, ”dzp”, ”$c$-pvtz”, ”tzp”. This calculation shows that any asymetry, originating from the presence of different atoms or here simply on the different treatment of the same atom, leads to the formation of a large electric dipole in the energy spectrum.
I.2.4 Choice of ansatz
A crucial piece of the VQE algorithm is the choice of the variational ansatz used in the calculation: one aims at a quantum circuit as shallow as possible, yet the ansatz must be sufficiently expressive to capture the ground state of the molecule and also permit the associated optimization. Of interest to us here is the scaling of the number of gates $N_{g}$ of the ansatz with the size $N$ of the molecule. A lower bound for $N_{g}$ is actually the number of free parameters of the ansatz since one gate can encode at most $O(1)$ parameters (in practice there are often many more gates than parameters, this is really a lower bound).
Following the difficulties encountered with the ”hardware efficient” ansatz (with respect in particular to the presence of barren plateaus), the community started to consider ansätze inspired by successful approaches in computational chemistry such as the UCC ansatz. These ansätze have relatively stiff scalings (e.g., $N^{6}$ for UCCSDT with single double and triple excitations, not taking into account overheads stemming from fermion-to-qubit encodings or energy measurements). However these are the scalings that are also needed in classical coupled cluster calculations, so it is very possible that this scaling is already close to optimal. Recently, some authors have attempted to use simpler ansätze. For instance, ADAPT-VQE [26] is inspired by a trotterized version of UCC and constructs the ansatz iteratively by adding the operators one by one from a pool of possible excitations. Due to the trotterization (often found in UCC-VQE too), the ansatz is somewhat closer to configuration interaction than coupled cluster and as such it is not clear how it performs with respect to size consistency. A significant reduction up to 60% was observed in the number of parameters used with respect to UCC with a better precision, at the cost of extra sampling. Other algorithms try to avoid the need for the optimization part (and associated gradient calculation) of VQE and try to find schemes that converge naturally to the ground state. One such construction approaches the ground state adiabatically through a randomized algorithm [38] that must be averaged over different circuits with random insertions. One could also construct a Lindbladian dynamics designed to converge to the ground state [39], use auxiliary qubits as a heat sink to cool down the system to the ground state [40], and engineer a transformed Hamiltonian [41] designed for cooling. Similar techniques can be used to compute Green’s functions in impurity models [42]. These techniques have mostly been demonstrated on very small scales. More research will be needed to study their potential for providing an advantage both at the practical and even theoretical level [43].
In any case, irrespectively of the type of ansatz and scheme used to optimize it towards the ground state, it is difficult to imagine that the scaling of $N_{g}$ would be more favourable than $M^{4}$. Indeed, one needs at the very least as many gates (hence angles) as the number of parameters that define the problem itself (the $M^{4}$ Coulomb integrals in (1)). What is more, such an optimum situation assumes that one has a perfect parametrization of the corresponding $M^{4}$-dimensional manifold inside the much larger $2^{M}$-dimensional Hilbert space, a very unlikely situation. Note a recent proposal [44] for approximate simulations of the dynamics where the scaling is reduced from $M^{4}$ for a Trotter simulation to $M^{2}$ at the cost of introducing extra qubits to rewrite the Coulomb tensor in a diagonal form. But whether this work can be extended to change the scaling of VQE ansatze remains unclear.
I.2.5 Application: criterion estimation for Benzene
@@
We are now in position to answer the question: what does the criterion Eq. (8) imply in practice? It is generally accepted that a quantum chemistry calculation must reach ”chemical accuracy”, $\eta_{\rm chem}=1{\rm kcal/mol}\ \approx 1.6{\rm mHa}\approx 500K$} (for energy differences). As we have seen, a very conservative estimate for the noise energy is $E_{\rm noise}\geq 1{\rm Ha}$, likely much more. For the $H_{2}$ molecule in a minimum basis set of just two orbitals per atom (STO-3G), one gets $E_{\infty}-E_{0}=1.02{\rm Ha}$} for the depolarizing noise, so that we estimate $E_{noise}-E_{V}$} to be of the order of 11Ha in this simple case. On the other hand, one needs a variational ansatz expressive enough to reach $\eta_{\rm chem}$, which provides a constraint $N_{g}\geq N_{g}(\eta_{\rm chem},N)$.
@@
To be concrete, we consider a recent blind test benchmark on the benzene molecule [45]. Benzene is a nontrivial calculation for classical approaches. Yet [45] showcased that a variety of classical techniques arrived at chemical precision using $30$ electrons distributed on $108$ orbitals. It can be considered as a minimum target for VQE: anything smaller is likely to be addressable by classical means. Using the UCC ansatz, inspired by the successful coupled cluster approach used in quantum chemistry, would require to include at least single, double and triple excitations (actually quadruple would probably be needed too), which translates into $N_{g}>N^{6}$ gates. This is probably a very conservative estimate that ignores the additional cost to implement the (non-local) fermionic operators with qubits as well as many prefactors. One arrives at a noise level $\epsilon\leq 10^{-3}/(30)^{6}$, which translates into $\epsilon\leq 10^{-12}$, that is many orders of magnitude below the best existing quantum hardware. Even if error mitigation could be maintained at a level of $G=0.001$ (which is doubtful for such a large system, see the discussion in the next paragraph), the corresponding bare error level needed would remain extremely small $\epsilon_{\rm bare}\leq 10^{-9}$ for a problem that is already within reach of several classical techniques.
@@
Such a low level of noise is difficult to imagine without a full-fledged implementation of quantum error correction. Since quantum error correction itself comes with an important overhead in term of computational time, one wants to avoid the accumulation of statistics required in VQE. We will thus turn to a different algorithm, QPE, in Sec. II.
I.2.6 Prospects of error mitigation
As we have argued in the above sections, decoherence introduces a large bias to the sought-after energy and is therefore very harmful. The reason for this bias being large is twofold: noise generically produces high-energy states, and is a multiplicative process (hence grows exponentially quickly). In this paragraph, we quickly discuss the possibility to use error mitigation techniques to alleviate this problem. We emphasize that the noise estimates at which we have arrived are agnostic to the method used to obtain them: They are estimates of the effective noise levels per gate needed to be achieved, possibly after error mitigation or even error correction, not necessarily estimates of the bare error level on the physical qubits.
@@ I.2.6 Prospects of error mitigation
Error mitigation techniques have become increasingly popular over the last few years [46, 47, 48, 49, 50, 28, 51, 52, 53]. These techniques, including probabilistic error cancellation, zero-noise extrapolation, assignment error mitigation, etc., essentially consist in reducing the bias on the energy estimator using more or less precise information on the noise processes at play. This comes at a steep price: the variance of the energy estimator increases exponentially with the number of gates of the circuit, and even, in some cases, the number of qubits [31, 54, 55]. This in turn, leads to a number of required shots, and thus a run time, that increase exponentially.
Let us be more specific with the concrete example of probabilistic error cancellation [47, 48]. In this method, to suppress the bias in energy in a noisy VQE, caused by the deviation between the noisy energy $E_{\mathrm{noise}}=\mathrm{Tr}(\rho_{\mathrm{noise}}H)$ and the perfect energy $E_{V}=\langle\Psi_{V}|H|\Psi_{V}\rangle$, one decomposes the “perfect” channels $\mathcal{E}_{\mathrm{perfect}}^{(k)}(\rho)=U^{(k)}\rho(U^{(k)})^{\dagger}$ (with $U^{(k)}$ the $k$th gate in the variational circuit) as a linear combination of the actual (noisy) channels $\mathcal{E}_{l}(\rho)$ that are realized in the quantum computer:
$\mathcal{E}_{\mathrm{perfect}}^{(k)}=\sum_{l}q_{l}^{(k)}\mathcal{E}_{l},$ (16)
with the underlying assumption that these noisy channels form an independent family, and the $\sum_{l}q_{l}^{(k)}=1$ to preserve the trace-preserving character (but the $q_{l}^{(k)}$ may be negative). The perfect energy can thus be decomposed as a very large sum, which is sampled through Monte Carlo:
$E_{\mathrm{perfect}}$ $=\sum_{l_{1}\dots l_{N_{\mathrm{g}}}}q_{l_{1}}^{(1)}\cdots q_{l_{N_{\mathrm{g}}}}^{(N_{\mathrm{g}})}\mathrm{Tr}\left(H\mathcal{E}_{l_{N_{\mathrm{g}}}}\circ\mathcal{E}_{l_{N_{\mathrm{g}}-1}}\circ\cdots\mathcal{E}_{l_{1}}(\rho_{\mathrm{i}})\right)$
$=\Gamma\sum_{l_{1}\dots l_{N_{\mathrm{g}}}}p_{l_{1}}^{(1)}\cdots p_{l_{N_{\mathrm{g}}}}^{(N_{\mathrm{g}})}\mathrm{s}(l_{1},\dots l_{N_{\mathrm{g}}})\mathrm{Tr}\left(H\mathcal{E}_{l_{N_{\mathrm{g}}}}\circ\mathcal{E}_{l_{N_{\mathrm{g}}-1}}\circ\cdots\mathcal{E}_{l_{1}}(\rho_{\mathrm{i}})\right)$ (17)
$\approx\frac{\Gamma}{N_{s}}\sum_{i=1}^{N_{s}}\mathrm{s}(l_{1}^{(i)},\dots l_{N_{\mathrm{g}}}^{(i)})\mathrm{Tr}\left(H\mathcal{E}_{l_{N_{\mathrm{g}}}^{(i)}}\circ\mathcal{E}_{l_{N_{\mathrm{g}}-1}^{(i)}}\circ\cdots\mathcal{E}_{l_{1}^{(i)}}(\rho_{\mathrm{i}})\right).$ (18)
Here, the probabilities $p_{l}^{(k)}=q_{l}^{(k)}/\sum_{l^{\prime}}q_{l^{\prime}}^{(k)}$ are introduced to deal with the possible negativity of the $q_{l}^{(k)}$’s, $\Gamma=\prod_{k=1}^{N_{g}}\sum_{l^{\prime}}\left|q_{l^{\prime}}^{(k)}\right|$ and $\mathrm{s}(l_{1},\dots l_{N_{\mathrm{g}}})=\prod_{k=1}^{N_{g}}\mathrm{sign}(q_{l_{k}}^{(k)})$. The advantage of the summand of Eq. (18) is that it can be estimated with the noisy computer at hand by running a circuit with (noisy) gates $\mathcal{E}_{l_{1}^{(i)}},\dots,\mathcal{E}_{l_{N_{\mathrm{g}}}^{(i)}}$ and measuring $H$ at the end, resulting in an unbiased estimate of $E_{\mathrm{perfect}}$. The price to pay is, however, twofold. First, one needs to know precisely the noise models of the hardware to be able to perform the decomposition of Eq. (16). This will require one to perform expensive tomography experiments, the cost of which will increase algebraically as one requires more precision or exponentially as one includes more complex noise processes with more qubits (such as crosstalk between different gates). There are also underlying assumptions such as the absence of drift (due to e.g. the two-level systems that are currently a very important limiting factor of superconducting qubits) or the precision of the tomography that ultimately limits the precision reachable by this approach. Second, and more fundamentally, the estimator comes with a statistical uncertainty $\Delta E_{\mathrm{shot}}$ that scales with the $\Gamma$ factor, which can be rewritten as $\Gamma=\prod_{k=1}^{N_{g}}\left(1+2\eta_{k}\right)$, with the so-called “negativity” $\eta_{k}=\sum_{l,q_{l}<0}q_{l}^{(k)}$. To get an intuition how it scales, let us suppose that all gates have the same negativity $\eta$. Then
$\Gamma\approx e^{2\eta N_{\mathrm{g}}}.$ (19)
Let us make a connection to our error threshold $\epsilon_{\mathrm{max}}$ (Eq. (8)). So far, our main concern was the bias $\Delta E$ in the energy, $\Delta E=\left(1-F\right)(E_{\mathrm{noise}}-E_{V})\approx\epsilon N_{g}(E_{\mathrm{noise}}-E_{V})$. For a finite number of shots, however, the total energy error has a contribution from this bias and from statistical shot noise:
$\Delta E_{\mathrm{tot}}=\sqrt{\Delta E^{2}+\Delta E_{\mathrm{shot}}^{2}}.$ (20)
With probabilistic error cancellation, we can a priori send $\Delta E$ to zero, but $\Delta E_{\mathrm{shot}}$ is multiplied by $e^{2\eta N_{g}}$. To keep $\Delta E_{\mathrm{shot}}$ constant, we thus need to increase $N_{\mathrm{s}}$ by a factor $e^{4\eta N_{g}}$.
Note that this exponential increase in the complexity is rather general: the exponential loss of fidelity can be viewed as a loss of information on the system. To fight it, one must acquire extra information either during the performance of the circuit (that would be quantum error correction with the syndrome measurements) or after (this is mitigation). In the latter case, the loss of information has had the time to accumulate exponentially and must be matched by some sorts of extra measurements. This argument applies irrespectively of the type of mitigation that is being used. The noise can only be mitigated to the extent that it is well understood. For instance, mitigating crosstalk proves to be much harder than simple one-qubit gate errors [56].
New classes of methods such as Refs. [57, 58] are delegating a much larger share of the calculation to classical supercomputers. In Refs. [57, 58] the role of the quantum computer is much narrower than in a VQE calculation: it merely proposes the relevant Slater determinants that will be included in a subsequent (purely classical) configuration-interaction type of calculation. Establishing the advantage of such techniques will require more studies. In particular, they face the same limitation as the classical techniques they are based on, so they can at best provide a limited improvement compared to those techniques. Should they provide a genuine improvement, it would also be interesting to study if the proposal step could be done classically.
We therefore think it is unlikely that error mitigation will solve the noise problem discussed above: in practice, the number of shots cannot be scaled exponentially with system size, leading to a residual bias. In any case, the above criterion still holds in the presence of practical error mitigation: One just needs to replace $\epsilon$ by $\epsilon^{\mathrm{\rm mitigated}}$ in Eq. (8). For small molecules with less than $15$ orbitals, it has been estimated [59] that error mitigation might improve by one or two orders of magnitude the maximum allowed error per gate for VQE (see also [52] for a 20-qubit case). For larger molecules, because of the scaling problems mentioned above, the gain one can expect from mitigation should be even lower, leaving the threshold of $10^{-12}$ for Benzene out of reach.
I.3 VQE versus its classical competitor, the variational Monte Carlo (VMC) algorithm
@@
(III) Hardware availability Need ”fault-tolerant quantum computer” level of precision (this work) Works on existing CPU, GPU and TPU. Trivial parallelism.
@@
Table 1: Comparison between VQE and VMC for the three main difficulties faced by these algorithms.
In the previous section, we discussed a necessary condition for VQE to be able to reach chemical accuracy. Here, we stress the fact that a condition for quantum advantage is more stringent: VQE must be faster than competing purely classical algorithms. There are essentially three classes of classical algorithms that we could think of. (i) The first class corresponds to algorithms designed to simulate quantum circuits such as in Refs. [60, 29, 61]. Trivially, if the quantum circuit of the VQE ansatz can be simulated, there is not much point in building the corresponding quantum computer. (ii) Second are state-of-the-art computational chemistry techniques. They involve a variety of approaches depending on the size of the molecules being studied, the accuracy needed and the type of constituents (e.g. organic molecules with close shells versus the presence of transition metals that imply the need for multi-reference calculations). These techniques include density functional theory, coupled cluster methods, various variants of configuration interaction, quantum Monte Carlo and the density matrix renormalisation group (DMRG) tensor network approach. All these techniques beat VQE working on existing quantum hardware by a large margin. The question is therefore not how they perform but how they scale when going to more complex and/or more accurate calculations. To perform a meaningful comparison with VQE in this context, we therefore consider (iii) a third class of classical techniques: approaches that share as many common points with VQE as possible, in particular the property of being stochastic.
A natural comparison point for VQE is the classical algorithm of variational Monte Carlo (VMC) [62]—an old technique that has been experiencing an important revival in the last few years thanks to the success of deep neural networks based ansatz. Indeed, the introduction of neural network variational wave functions with a large number of parameters $N_{\theta}$ has led to very encouraging results [63, 64]. In VMC, one also considers a variational wavefunction $|\Psi_{V}(\vec{\theta})\rangle$. This function can be anything as long as one is able to compute its value $\langle\vec{s}|\Psi_{V}\rangle$ for a given configuration $|\vec{s}\rangle$. This includes wavefunctions explicitly written in term of a neural network as well as physically-motivated wavefunctions, or combinations thereof. However, these ansätze are typically not written in terms of a quantum circuit in order to avoid the large computational overhead associated with the contraction of the associated tensor network. In VMC, the variational energy is expressed as
@@
The sum over configurations in (21) is sampled by a Monte Carlo algorithm, namely one draws $N_{s}$ samples from the distribution $|\langle\vec{s}|\Psi_{V}\rangle|^{2}$ and estimates the energy $E_{V}$ as
@@
Here, even if $|\Psi_{V}\rangle$ happens to be the exact ground state, it is not an eigenstate of the individual $P_{\alpha}$ operators and the variance thus does not vanish. In fact, since those terms originally stem from the kinetic and Coulomb repulsion matrix elements, they have quite large magnitude (typically in the $1$ Ha range). This absence of variance reduction in VQE reflects the fact that it is much more advantageous to measure the full Hamiltonian at once than its different components separately. In practice, as the optimization proceeds in VMC, one wants to measure $E_{V}$ with more and more digits of precision. Since $|\Psi_{V}\rangle$ gets closer to $|\Psi_{0}\rangle$, in VMC the energy variance decreases and hence the precision of the calculation also becomes better for the same experimental time (same number of samples). In contrast, in VQE the variance of the different energies $E_{\alpha}$ does not decrease, hence one must increase the number of samples by a factor 100 for each new digit in precision on the measure of $E_{V}$. This might constitute another important advantage in favor of VMC.
@@
(5) The two techniques have access to different sorts of ansätze. VMC is more versatile in that respect and can simulate VQE ansätze (as long as the quantum circuit depth is not too deep) while the reverse is not true. For instance, VMC can use Tensor network states that give an efficient representation of weakly entangled states. But VMC can also use neural network states which can sustain highly entangled states including volume law entanglement. In VQE, the quantum nature of the hardware allows, in principle, to build arbitrarily entangled states, but in practice the depth might be excessive. This is particularly true for the existing (NISQ) hardware which is limited to shallow circuits, hence limited entanglement. At the moment it is unclear if one technique has an advantage in that respect. One should keep in mind that the expressivity of the ansatz is strongly linked to the ease/difficulty to perform the optimization. An ansatz which is too expressive (for instance the so-called ”hardware efficient ansatz” in VQE) leads to barren plateaus as most of the parameter space corresponds to almost random (chaotic) states (see the above discussion on barren plateaus).
The conclusions of this comparison between VQE and VMC are summarized in Table 1. What one should take away from this exercise is that, in contrast to e.g. Shor’s algorithm, which is expected to speed up the factorization of large numbers exponentially, there are no similar indications for VQE, quite the opposite actually. Even if we had perfect quantum computers, VMC might still remain superior to VQE. (See also [66] for a discussion including other classical Monte Carlo methods).